OpenStudy (anonymous):

What is the radius of the incircle of the triangle whose sides measure 5, 12 and 13 units?

OpenStudy (triciaal):

do you recognize this is a right angle triangle?

OpenStudy (mathstudent55):



OpenStudy (anonymous):

AH! yeah i get it now. I know radius of in circle is given by R=2a/p a= area and p= perimeter..... so i need to calculate perimeter as well as area of this Rt. angled triangle. Thanks @triciaal @mathstudent55

OpenStudy (anonymous):

True also u can also see that in a right angled triangle, the radius of the incircle is given by the following relation \[\frac{ \sum od perpendicular sides-hypotenuse }{ 2 }\] @pathfinder_007

OpenStudy (anonymous):

yes yes i will get 5+12-13/2= 2 units(answer) thanks @aryandecoolest